# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v3

@@ III.3 Fronts and localized structures
In a one-dimensional space, fronts manifest as heteroclinic orbits that bridge two distinct uniform states, potentially differing in stability Homburg and Sandstede (2010). These dynamics give rise to two front types, distinguished by their approach to the uniform state, specifically the nature of their ’tails’. Tails that approach uniformly without oscillations form what are known as tame or flat fronts. Conversely, tails that exhibit damped oscillations around the uniform state result in wild fronts. Illustrations of both front types within the FHN framework are provided in McKean Jr (1970); Zemskov et al. (2008, 2015) and are depicted in Fig. 10A.
Near uniform states, front behaviors can be linearly approximated by:
$\begin{pmatrix}u(x)\\ v(x)\end{pmatrix}-\begin{pmatrix}u_{h}\\ v_{h}\end{pmatrix}\propto e^{\lambda x},$ (26)
@@
Like all nonlinear phenomena, front dynamics can exhibit instabilities, leading to the formation of complex patterns Hagberg and Meron (1994a, b). Specifically, fronts can become susceptible to transverse modulations, a phenomenon known as transverse front instability, resulting in the ”fingering” effect and the subsequent development of ”labyrinth patterns” Hagberg et al. (2006). Additionally, fronts might experience the nonequilibrium Ising-Bloch instability, giving rise to two counterpropagating fronts Hagberg and Meron (1994b). However, in the context of the FHN model, the coexistence of a stable planar front with a stable large-amplitude stripe pattern, a prerequisite for nonlinear front transverse instability, has yet to be observed Hagberg et al. (2006).
McKean, in 1969, introduced a simplified approach to the FHN model through a piecewise approximation (Fig. 10C), facilitated by the distinct time scales in the system McKean Jr (1970). This approximation transforms the cubic function into a piecewise or Heaviside function, allowing for an analysis of front formation and propagation within an integro-differential framework of the FHN model:
@@
Stable fronts result in regular wave propagation, contingent on the presence of an ”energetically preferred” state. This is illustrated in scenarios where regions of high and low activity eventually converge to a preferred state, as shown in Fig. 10B. The concept of ”equivalent uniform” states, where two states are related by the $u\rightarrow-u$ transformation and possess equal energy in systems with definable free energy, leads to static fronts Chomaz (1992). This equilibrium is attained at the Maxwell point of the system, identified in the FHN model at $a=a_{M}\equiv 0$. Front movement is observed when parameters deviate from this point.
@@
$\frac{dD}{dt}=Ae^{{\rm Re}(\lambda)D}{\rm cos}({\rm Im}(\lambda)D)+B,$ (28)
@@
In contrast, when the fronts are wild (${\rm Im}(\lambda)\neq 0$), any stationary solution of Eq. (28) corresponds to the locking or pinning of two fronts, resulting in localized states. Near the system’s Maxwell point ($a=a_{M}$, where $B=0$) in an infinite domain, Eq. (28) presents countless equilibrium distances with alternating stability Parra-Rivas et al. (2023b). Deviating from the Maxwell point ($B\neq 0$), the variety of equilibrium distances—and consequently, the potential localized states (LSs)—decreases. The emergence of two distinct LSs near the Maxwell point is depicted in Fig. 10E for $a=10^{-5}$ (left) and $a=10^{-3}$ (right). In the former scenario, the fronts lock very fast, resulting in a broad LS, while for $a=10^{-3}$, the slower convergence of fronts eventually leads to the formation of a narrower, single-peak LS.
Figure 11: Excitable traveling pulses in the FHN model. A. Formation process of an excitable traveling pulse. Initially, the system is disrupted beyond a certain pseudothreshold determined by the unique structure of the folded nullclines. This leads to the diffusion-induced excitation of neighboring regions, causing the pulse to travel through space. Subsequently, areas that were excited earlier trace back along the nullcline, eventually settling into a uniform state. B. A simplified representation of the nullclines as piecewise linear functions, an approximation that holds primarily under significant time scale separation. C. Two types of pulses: a ’tame’ pulse (left), generated under the parameters $(D_{u},D_{v},a,b,\varepsilon)=(1,1,0.2,2,0.01)$, and a ’wild’ pulse (right), arising from $(1.5,1,0.4,0.5,0.1)$. D. Propagation dynamics of the ’wild’ pulse. The pulse maintains its shape as it moves, shown at two distinct moments, $t_{0}$ and $t_{1}$, marked by vertical dashed lines on the kymograph.
@@ III.4 Traveling pulses
[…]
@@
While the generation of excitable traveling pulses is conceptually rather simple, their investigation is complex. The FHN model is extensively employed to explore key features of traveling pulses, utilizing its piecewise linear approximation for simplicity (refer to Fig. 11B). The insights from earlier discussions apply here, demonstrating the broader relevance of the piecewise linear FHN model beyond McKean’s work McKean Jr (1970), with various studies addressing traveling pulses Majumdar and Lahiri (2007); Rinzel and Keller (1973); Zemskov et al. (2018); Zemskov and Epstein (2010) and traveling waves Zemskov et al. (2015); Rotstein et al. (2012) within this framework.
Revisiting the categorization from Section III.3, traveling pulses can also be distinguished as tame or wild based on their approach to the uniform state. Fig. 11C.1 and C.2 show examples of tame (with parameters $D_{u}=1$, $D_{v}=1$, $a=0.2$, $b=2$, $\varepsilon=0.01$) and wild (with parameters $D_{u}=1.5$, $D_{v}=1$, $a=0.4$, $b=0.5$, $\varepsilon=0.1$) pulses, respectively. The propagation of a wild pulse is depicted in Fig. 11D, where snapshots at times $t_{0}$ and $t_{1}$—marked by white dashed lines in the kymograph—illustrate the pulse’s consistent shape throughout its journey.
@@ III.5 Pacemakers and wave trains
We now look into the dynamics of oscillatory media coupled via diffusion, a subject less frequently studied in existing research on the FHN model Stich et al. (2009); Rabinovitch et al. (2001); Elphick et al. (1990); Rombouts and Gelens (2020, 2021). This aspect is particularly relevant to cellular and molecular biology, where traveling waves facilitate long-distance system synchronization and intra- or inter-cellular communication Gelens et al. (2014); Nolet et al. (2020a); Beta and Kruse (2017); Di Talia and Vergassola (2022). For these biological processes to coordinate effectively, it is crucial for the waves to propagate sufficiently quickly through the medium.
@@
Mathematically, wave train studies transcend the FHN context Hagan (1981); Dockery et al. (1988) due to their widespread applicability. Utilizing the FHN model to describe wave trains, regularly used in cell cycle contexts Rombouts and Gelens (2020); Nolet et al. (2020b); Di Talia and Vergassola (2022); Gelens et al. (2014); Nolet et al. (2020a), underscores its capacity to mimic target patterns resembling the behavior of a nucleus or even multiple nuclei within the cellular context Nolet et al. (2020a, b). Such patterns are not unique to the FHN model but are also seen in chemical Taylor (2002), cardiac Alonso et al. (2016), and neuronal models Muller et al. (2018), underscoring the FHN’s conceptual utility in exploring relaxation dynamics across various scientific domains.
Figure 14: Dynamics of 2D spatio-temporal patterns. A. The dynamic interplay between the homogeneous state’s oscillatory dynamics and spatial pattern is shown, set within the Turing-Hopf regime for parameters $(a,\varepsilon,D_{u},D_{v})=(0.025,0.5,1,5)$. In the upper series ($b=1.1$), a striped pattern emerges, generated cycle-by-cycle through periodic oscillations, which progressively takes over the oscillatory domain, leading to a fully established striped pattern across the system. Conversely, the lower series ($b=0.8$) shows a distinct dynamic equilibrium, where the system bifurcates into oscillatory and stationary patterned halves. At the interface, oscillatory dynamics verge on stabilization, while the adjacent striped pattern exhibits rhythmic fluctuations, akin to a ’breathing stripe’, highlighting a unique boundary behavior. B. Spiral wave formation, originating from a phase defect induced by gradient-like initial conditions for $u$ (horizontal) and $v$ (vertical), with parameters $(a,b,\varepsilon,D_{u},D_{v})=(0.1,1,0.005,1,1)$ (oscillatory regime). A kymograph, representing a horizontal cross-section through the spiral’s core, captures the phase defect. The numerical simulations in A and B use Neumann boundary conditions.
III.6 Complex spatiotemporal dynamics in two spatial dimensions
@@ III.6.1 Universal phenomena of pattern formation
The exploration of the bifurcation structure of localized states and patterns, crucial for understanding a system’s self-organizational response, has been thorough in various spatially extended models but remains incomplete for the FHN model. In biology, for instance, the morphogenetic patterns have been examined using the Gierer–Meinhardt model Al Saadi and Champneys (2021), while chemical pattern formations have been delineated through the Gray-Scott Al Saadi and Champneys (2021) and Brusselator Uecker and Wetzel (2020) models. Ecological studies have extensively investigated vegetation patterns to pinpoint tipping points critical for understanding phenomena like desertification Gowda et al. (2014); Gandhi et al. (2018); Zelnik et al. (2018, 2013); Meron et al. (2019); Al Saadi and Parra-Rivas (2023); Parra-Rivas and Fernandez-Oto (2020); Ruiz-Reynés et al. (2020a, b).
@@ III.6.2 Stationary patterns in two spatial dimensions
In our exploration of the FHN model, we turn our attention to two spatial dimensions to unveil the bifurcation diagram of spatial structures. While our focus is on two dimensions, it is important to note that research has extended into higher-dimensional pattern investigations, offering a richer understanding of spatial dynamics Li et al. (2022); Zaritski et al. (2004).
Stationary patterns, stemming from Turing bifurcations, have been a significant area of study. Various researchers have looked into the genesis of these spatial formations, formulating amplitude equations Iqbal et al. (2017); Zheng and Shen (2015). A notable challenge is the close proximity of Turing and Hopf bifurcations, complicating pattern identification. To circumvent this, strategies involve operating near the Turing threshold or employing white noise to steer the system towards stable pattern branches. Our numerical simulations reveal a diversity of stable spatial patterns across different control parameters, specifically for $b=1.26$, $\varepsilon=0.5$, $D_{u}=1$, and $D_{v}=5$, as illustrated in Fig. 13. These patterns, transitioning from negative to positive spots through labyrinthine structures, underscore the labyrinth configuration’s robust stability for the chosen parameters. Notably, Goldstein’s work on interface growth within the FHN framework Goldstein et al. (1996), inspired by chemical front interactions Lee et al. (1993), aligns with our findings in the labyrinth domains (Fig. 13B).
@@ III.6.4 The Turing-Hopf regime
The interplay between oscillatory dynamics and pattern formation in the Turing-Hopf regime introduces a diverse array of spatio-temporal patterns, including expanding rings, pulsating waves, oscillatory patterns, and self-replicating spots Hagberg and Meron (1994a); Vasiev (2004); Li et al. (2022). This regime, primarily conceptualized through the Ginzburg-Landau equation Walgraef (2012), remains less explored in the FHN model, presenting an opportunity to deepen our understanding of systems where pattern formation and oscillatory behaviors coexist Hagberg and Meron (1994a); Brauns et al. (2021); Lee et al. (1994); Lee (1997); Hladyshau et al. (2021). Notably, phenomena such as ’breathing stripes’ exemplify the complex dynamics within this regime, as illustrated in Fig. 14A. Here, we show two scenarios: one where an initial pattern progressively dominates the oscillatory landscape and another featuring a dynamic equilibrium between oscillatory and patterned regions, highlighted by the breathing stripe at the boundary.
@@ III.6.5 Spiral waves
For the latter, we focus on spirals, which have been widely studied in the FHN model for mathematical purposes Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993a); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) and for their applicability in cardiology Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), where spiral behavior has been associated with arrhythmias. From the simple spirals studied by Erhaud and Winfree Meron (1992); Winfree (1991b) to the complex collisions studied by the latter Winfree (1991a), and through spiral turbulence emerging from the loss of stability of these structures Zaritski et al. (2004); Bär and Eiswirth (1993a), the FHN model has been utilized to characterize several regimes associated with spiral behavior. In Fig. 14B, we illustrate the creation of a stable spiral arising from a phase defect induced by considering horizontal gradient-like initial conditions for $u$ and horizontal gradient-like initial conditions for $v$ in an oscillatory regime ($a=0.1$, $b=1$, $\varepsilon=0.005$, $D_{u}=1$, $D_{v}=1$).
Spiral wave patterns, particularly relevant in cardiological contexts due to their association with arrhythmic behaviors Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), represent another facet of the FHN model’s versatility. From the foundational works on simple spirals Meron (1992); Winfree (1991b) to the studies on spiral turbulence Zaritski et al. (2004); Bär and Eiswirth (1993a) and collision dynamics Winfree (1991a), the model serves as a vital tool in understanding the nuances of spiral behavior. Fig. 14B demonstrates the formation of a stable spiral, a phenomenon induced by specific initial conditions (here a phase defect) within an oscillatory setting. This insight into spiral formation within the FHN model not only advances mathematical explorations Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993a); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) but also has profound implications in the study of cardiac arrhythmias Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), offering potential avenues for therapeutic interventions.

### v4

@@ III.3 Fronts and localized structures
Figure 10: Dynamics of localized states through the lens of various front behaviors. A. Two contrasting front types - a “tame” front (left) and a “wild” front (right), with the tame front parameters set at $(D_{u},D_{v},a,b,\varepsilon)=(1,1,0,2,0.01)$ and the wild front at $(0.49,1,0.01,1.1,1)$. B. Sketch of how a traveling front can be created. C. The commonly used piecewise linear approximation leverages the separation of time scales to analytically predict front properties. D. Kymographs and their corresponding profiles for traveling tame fronts, employing the same parameters as the tame front with added asymmetry by adjusting $a$ to 0.1 (left) and -0.1 (right). E. Formation of localized states via front interactions, with $a$ values set at 10-5 (left) and 10-3 (right) for the parameters $(D_{u},D_{v},b,\varepsilon)=(0.49,1,1.1,1)$.
In a one-dimensional space, fronts manifest as heteroclinic orbits that bridge two distinct uniform states, potentially differing in stability Homburg and Sandstede (2010). These dynamics give rise to two front types, distinguished by their approach to the uniform state, specifically the nature of their ‘tails’. Tails that approach uniformly without oscillations form what are known as tame or flat fronts. Conversely, tails that exhibit damped oscillations around the uniform state result in wild fronts. Illustrations of both front types within the FHN framework are provided in McKean (1970); Zemskov et al. (2008, 2015) and are depicted in Fig. 10A.
Near homogenous states, front behaviors can be linearly approximated by:
$\begin{pmatrix}u(x)\\ v(x)\end{pmatrix}-\begin{pmatrix}u_{\text{HS}}\\ v_{\text{HS}}\end{pmatrix}\propto e^{\lambda x},$ (25)
@@
Like all nonlinear phenomena, front dynamics can exhibit instabilities, leading to the formation of complex patterns Hagberg and Meron (1994a, b). Specifically, fronts can become susceptible to transverse modulations, a phenomenon known as transverse front instability, resulting in the “fingering” effect and the subsequent development of “labyrinth patterns” Hagberg et al. (2006). Additionally, fronts might experience the nonequilibrium Ising-Bloch instability, giving rise to two counterpropagating fronts Hagberg and Meron (1994b). However, in the context of the FHN model, the coexistence of a stable planar front with a stable large-amplitude stripe pattern, a prerequisite for nonlinear front transverse instability, has yet to be observed Hagberg et al. (2006).
McKean, in 1969, introduced a simplified approach to the FHN model through a piecewise approximation (Fig. 10C), facilitated by the distinct time scales in the system McKean (1970). This approximation transforms the cubic function into a piecewise or Heaviside function, allowing for an analysis of front formation and propagation within an integro-differential framework of the FHN model:
@@
Stable fronts result in regular wave propagation, contingent on the presence of an “energetically preferred” state. This is illustrated in scenarios where regions of high and low activity eventually converge to a preferred state, as shown in Fig. 10B. The concept of “equivalent uniform” states, where two states are related by the $u\rightarrow-u$ transformation and possess equal energy in systems with definable free energy, leads to static fronts Chomaz (1992). This equilibrium is attained at the Maxwell point of the system, identified in the FHN model at $a=a_{M}\equiv 0$. Front movement is observed when parameters deviate from this point.
@@
$\frac{dD}{dt}=Ae^{{\rm Re}(\lambda)D}\cos({\rm Im}(\lambda)D)+B,$ (27)
@@
In contrast, when the fronts are wild (${\rm Im}(\lambda)\neq 0$), any stationary solution of Eq. 27 corresponds to the locking or pinning of two fronts, resulting in localized states. Near the system’s Maxwell point ($a=a_{M}$, where $B=0$) in an infinite domain, Eq. 27 presents countless equilibrium distances with alternating stability Parra-Rivas et al. (2023b). Deviating from the Maxwell point ($B\neq 0$), the variety of equilibrium distances – and consequently, the potential localized states (LSs) – decreases. The emergence of two distinct LSs near the Maxwell point is depicted in Fig. 10E for $a=10^{-5}$ (left) and $a=10^{-3}$ (right). In the former scenario, the fronts lock very fast, resulting in a broad LS, while for $a=10^{-3}$, the slower convergence of fronts eventually leads to the formation of a narrower, single-peak LS Frohoff-Hülsmann et al. (2023).
@@ III.4 Traveling pulses
Figure 11: Excitable traveling pulses in the FHN model. A. Formation process of an excitable traveling pulse. Initially, the system is disrupted beyond a certain pseudothreshold determined by the unique structure of the folded nullclines. This leads to the diffusion-induced excitation of neighboring regions, causing the pulse to travel through space. Subsequently, areas that were excited earlier trace back along the nullcline, eventually settling into a uniform state. B. A simplified representation of the nullclines as piecewise linear functions, an approximation that holds primarily under significant time scale separation. C. Two types of pulses: a ‘tame’ pulse (left), generated under the parameters $(D_{u},D_{v},a,b,\varepsilon)=(1,1,0.2,2,0.01)$, and a ‘wild’ pulse (right), arising from $(1.5,1,0.4,0.5,0.1)$. D. Propagation dynamics of the ‘wild’ pulse. The pulse maintains its shape as it moves, shown at two distinct moments, $t_{0}$ and $t_{1}$, marked by vertical dashed lines on the kymograph.
@@
In one-dimensional systems, traveling pulses correspond to homoclinic orbits of the spatial system in the comoving reference frame. These orbits form connections from a stable fixed point back to itself Glendinning (1994). Illustrated in Fig. 11A, perturbations exceeding a critical threshold from a uniform state induce an increase in $u$, initiating a substantial excursion that, through diffusion, activates adjacent regions. The system eventually reverts to the initial state, descending along the folded nullcline before a sharp reduction in $u$ leads back to the baseline state along the opposing nullcline. This propagation mechanism is akin to a trigger wave initiating at the forefront, succeeded by a phase wave, a process underpinned by the excitable medium’s characteristics, facilitating the formation of traveling pulses Tyson and Keener (1988b).
@@
While the generation of excitable traveling pulses is conceptually rather simple, their investigation is complex. The FHN model is extensively employed to explore key features of traveling pulses, utilizing its piecewise linear approximation for simplicity (Fig. 11B). The insights from earlier discussions apply here, demonstrating the broader relevance of the piecewise linear FHN model beyond McKean’s work McKean (1970), with various studies addressing traveling pulses Majumdar and Lahiri (2007); Rinzel and Keller (1973); Zemskov et al. (2018); Zemskov and Epstein (2010) and traveling waves Zemskov et al. (2015); Rotstein et al. (2012) within this framework.
Revisiting the categorization from III.3, traveling pulses can also be distinguished as tame or wild based on their approach to the uniform state. Fig. 11C.1 and C.2 show examples of tame (with parameters $D_{u}=1$, $D_{v}=1$, $a=0.2$, $b=2$, $\varepsilon=0.01$) and wild (with parameters $D_{u}=1.5$, $D_{v}=1$, $a=0.4$, $b=0.5$, $\varepsilon=0.1$) pulses, respectively. The propagation of a wild pulse is depicted in Fig. 11D, where snapshots at times $t_{0}$ and $t_{1}$ – marked by white dashed lines in the kymograph – illustrate the pulse’s consistent shape.
@@ III.5 Pacemakers and wave trains
We now look into the dynamics of oscillatory media coupled via diffusion, a subject less frequently studied in existing research on the FHN model Stich et al. (2009); Rabinovitch et al. (2001); Elphick et al. (1990); Rombouts and Gelens (2020, 2021). This aspect is particularly relevant to cellular and molecular biology, where traveling waves facilitate long-distance system synchronization and intra- or inter-cellular communication Gelens et al. (2014); Beta and Kruse (2017); Di Talia and Vergassola (2022).For these biological processes to be coordinated effectively, it is crucial for the waves to propagate sufficiently quickly through the medium.
@@
Mathematically, wave train studies transcend the FHN context Hagan (1981); Dockery et al. (1988) due to their widespread applicability. Utilizing the FHN model to describe wave trains, regularly used in cell cycle contexts Rombouts and Gelens (2020); Nolet et al. (2020b); Di Talia and Vergassola (2022); Gelens et al. (2014); Nolet et al. (2020a); Cebrián-Lacasa et al. , underscores its capacity to mimic target patterns resembling the behavior of a nucleus or even multiple nuclei within the cellular context Nolet et al. (2020a, b); Cebrián-Lacasa et al. . Such patterns and their response to time-scale separation and to the presence of spatial heterogeneity are not unique to the FHN model. They are also seen in chemical Taylor (2002), cell cycle Chang and Ferrell Jr (2013); Puls* et al. (2024), cardiac Alonso et al. (2016), and neuronal models Muller et al. (2018), underscoring the FHN’s conceptual utility in exploring relaxation dynamics across various scientific domains, both theoretically and experimentally.
Figure 14: Dynamics of 2D spatio-temporal patterns. A. The dynamic interplay between the homogeneous state’s oscillatory dynamics and spatial pattern is shown, set within the Turing-Hopf regime for parameters $(a,\varepsilon,D_{u},D_{v})=(0.025,0.5,1,5)$. In the upper series ($b=1.1$), a striped pattern emerges, generated cycle-by-cycle through periodic oscillations, which progressively takes over the oscillatory domain, leading to a fully established striped pattern across the system. Conversely, the lower series ($b=0.8$) shows a distinct dynamic equilibrium, where the system bifurcates into oscillatory and stationary patterned halves. At the interface, oscillatory dynamics verge on stabilization, while the adjacent striped pattern exhibits rhythmic fluctuations, akin to a ‘breathing stripe’, highlighting a unique boundary behavior. B. Spiral wave formation, originating from a phase defect induced by gradient-like initial conditions for $u$ (horizontal) and $v$ (vertical), with parameters $(a,b,\varepsilon,D_{u},D_{v})=(0.1,1,0.005,1,1)$ (oscillatory regime). A kymograph, representing a horizontal cross-section through the spiral’s core, captures the phase defect. The numerical simulations in A and B use Neumann boundary conditions.
III.6 Complex spatio-temporal dynamics in two spatial dimensions
@@ III.6.1 Universal phenomena of pattern formation
The exploration of the bifurcation structure of localized states and patterns, crucial for understanding a system’s self-organizational response, has been thorough in various spatially extended models but remains incomplete for the FHN model. In biology, for instance, the morphogenetic patterns have been examined using the Gierer-Meinhardt model Al Saadi and Champneys (2021), while chemical pattern formations have been delineated through the Gray-Scott Al Saadi and Champneys (2021) and Brusselator Uecker and Wetzel (2020) models. Ecological studies have extensively investigated vegetation patterns to pinpoint tipping points critical for understanding phenomena like desertification Gowda et al. (2014); Gandhi et al. (2018); Zelnik et al. (2018, 2013); Meron et al. (2019); Al Saadi and Parra-Rivas (2023); Parra-Rivas and Fernandez-Oto (2020); Ruiz-Reynés et al. (2020a, b).
@@ III.6.2 Stationary patterns in two spatial dimensions
In our exploration of the FHN model, we turn our attention to two spatial dimensions to study the bifurcation diagram of spatial structures. While our focus is on two dimensions, it is important to note that research has extended into higher-dimensional pattern investigations, offering a richer understanding of spatial dynamics Li et al. (2022); Zaritski et al. (2004).
Stationary patterns, stemming from Turing bifurcations, have been a significant area of study. Various researchers have looked into the genesis of these spatial formations, formulating amplitude equations Iqbal et al. (2017); Zheng and Shen (2015). A notable challenge is the close proximity of Turing and Hopf bifurcations, complicating pattern identification. To circumvent this, strategies involve operating near the Turing threshold or employing white noise to steer the system towards stable pattern branches. Our numerical simulations reveal a diversity of stable spatial patterns across different control parameters, specifically for $b=1.26$, $\varepsilon=0.5$, $D_{u}=1$, and $D_{v}=5$, as illustrated in Fig. 13. These patterns, transitioning from negative to positive spots through labyrinthine structures, underscore the labyrinth configuration’s robust stability for the chosen parameters. Goldstein’s work on interface growth within the FHN framework Goldstein et al. (1996), inspired by chemical front interactions Lee et al. (1993), aligns with our findings in the labyrinth domains (Fig. 13B).
@@ III.6.4 The Turing-Hopf regime
The interplay between oscillatory dynamics and pattern formation in the Turing-Hopf regime introduces a diverse array of spatio-temporal patterns, including expanding rings, pulsating waves, oscillatory patterns, and self-replicating spots Hagberg and Meron (1994a); Vasiev (2004); Li et al. (2022). This regime, primarily conceptualized through the Ginzburg-Landau equation Walgraef (2012), remains less explored in the FHN model, presenting an opportunity to deepen our understanding of systems where pattern formation and oscillatory behaviors coexist Hagberg and Meron (1994a); Brauns et al. (2021); Lee et al. (1994); Lee (1997); Hladyshau et al. (2021). Phenomena such as ’breathing stripes’ exemplify the complex dynamics within this regime, as illustrated in Fig. 14A. Here, we show two scenarios: one where an initial pattern progressively dominates the oscillatory landscape and another featuring a dynamic equilibrium between oscillatory and patterned regions, highlighted by the breathing stripe at the boundary.
@@ III.6.5 Spiral waves
Spiral waves have been widely studied in the FHN model for mathematical purposes Ambrosio and Aziz-Alaoui (2012); Keener (1986b); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993b); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) and for their applicability in cardiology Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), where spiral behavior has been associated with arrhythmias. From the simple spirals studied by Erhaud and Winfree Meron (1992); Winfree (1991b) to the complex collisions studied by the latter Winfree (1991a), and through spiral turbulence emerging from the loss of stability of these structures Zaritski et al. (2004); Bär and Eiswirth (1993b), the FHN model has been utilized to characterize several regimes associated with spiral behavior. In Fig. 14B, we illustrate the creation of a stable spiral arising from a phase defect induced by considering horizontal gradient-like initial conditions for $u$ and horizontal gradient-like initial conditions for $v$ in an oscillatory regime ($a=0.1$, $b=1$, $\varepsilon=0.005$, $D_{u}=1$, $D_{v}=1$).
Spiral wave patterns, particularly relevant in cardiological contexts due to their association with arrhythmic behaviors Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), represent another facet of the FHN model’s versatility. From the foundational works on simple spirals Meron (1992); Winfree (1991b) to the studies on spiral turbulence Zaritski et al. (2004); Bär and Eiswirth (1993b) and collision dynamics Winfree (1991a), the model serves as a vital tool in understanding the nuances of spiral behavior. Fig. 14B demonstrates the formation of a stable spiral, a phenomenon induced by specific initial conditions (here a phase defect) within an oscillatory setting. This insight into spiral formation within the FHN model not only advances mathematical explorations Ambrosio and Aziz-Alaoui (2012); Keener (1986b); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993b); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) but also has profound implications in the study of cardiac arrhythmias Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008), offering potential avenues for therapeutic interventions.